### v1

@@ 4.4.2 Key propositions
By (24) in Corollary 3.9,
$\quad\,\mathbb{P}_{n}\Big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\,\Big|\,\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m})\text{ are distinct}\Big)$
$=1-\dfrac{m\sigma_{n}^{2}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$
Substituting into (78), we obtain
@@
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)$
$\quad\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
@@
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)$
$\quad\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
@@
This proves the lemma.
Proof 4.24 (Proof of Proposition 4.19 (2)).
@@
Fix vertices $v^{j}_{l},\,v^{j^{\prime}}_{l^{\prime}}$ such that $h_{r}<j\leq j^{\prime}\leq h_{r-1}$, $1\leq l\leq q_{n}(j)-m$, and $1\leq l^{\prime}\leq q_{n}(j^{\prime})-m$. It suffices to show that, uniformly for any $0\leq k<j-h_{r}$ and $1\leq k^{\prime}\leq j^{\prime}-h_{r}$,
${\mathbb{P}}_{n}\big(\tau(v^{j}_{l})=\infty~\big|~\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big)$ (79)
$=$ $(1+o_{n}(1))\prod_{h=1}^{j-(j^{\prime}-k^{\prime})-2}\Bigg[1-\frac{(m+1)\sigma^{2}}{q_{n}(\bar{h}+j-h)}(1+o_{n}(1))\Bigg]$
$\quad\times\prod_{h=j-(j^{\prime}-k^{\prime})}^{j-k}\Bigg[1-\frac{m\sigma^{2}}{q_{n}(\bar{h}+j-h)}(1+o_{n}(1))\Bigg].$
@@
$p_{n}(h):={\mathbb{P}}_{n}\big(\tau(v^{j}_{l})>h+1~\big|~\tau(v^{j}_{l})>h,\,\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big).$
Under this conditioning, $j^{\prime}-k^{\prime}$ represents the height at which $v^{j^{\prime}}_{l^{\prime}}$ merges into $U_{i,r}$. By exchangeability, taking $h_{r-1}$ as the base height, we have:
@@
    forms an (jh,m+2)(j-h,m+2)-uniform selection pair.
@@
    forms an (jh,m+1)(j-h,m+1)-uniform selection pair.
@@
    Uniformly for $0\leq h\leq j-h_{r}-1$ with $j-h>j^{\prime}-k^{\prime}+1$,
@@
    Uniformly for $0\leq h\leq j-h_{r}-1$ with $j-h\leq j^{\prime}-k^{\prime}$,
@@
    $q_{n}(h)$ $=\mathbb{P}_{n}\Big(\{\mathfrak{p}(u_{1}^{j-h})=\mathfrak{p}(u_{2}^{j-h})\}\cap\{\mathfrak{p}(u_{2}^{j-h}),\ldots,\mathfrak{p}(u_{m+1}^{j-h})\text{ are distinct}\}~\Big|~\text{same event}\Big)$
@@
This completes the proof.
@@
In the previous section we established the tightness of $\widetilde{\mathscr{T}}^{n}$. Here we first prove, under (H1), (H2’) and (H2”), the finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$ (Section 5.1); combined with tightness, this yields convergence of the contour functions. We then show in Section 5.2 that the discrepancy between the height and contour functions vanishes in the limit, so both processes converge to the same limiting process. Finally, we verify that (H2) implies (H2’) and (H2”) (Proposition 5.8), which completes the proof of Theorem 1.1.
@@ 5.1 Finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$
Recall that $\mathcal{T}^{\ell,\sigma}$ is the rooted ordered real tree encoded by $W^{\ell^{\sigma}}_{\alpha^{-1}(I(\ell)t)}$, and that $\mathcal{T}^{\ell,\sigma}_{k}$ denotes its uniform $k$-point subtree, whose law is described in Theorem 2.20. Also recall that $\widetilde{\mathscr{T}}^{n}$ is the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. Let $(V_{n,1},\dots,V_{n,\#V(\mathscr{T}^{n})})$ be a uniform random ordering of $V(\mathscr{T}^{n})$, and let $\widetilde{\mathscr{T}}^{n}_{k}$ be the rooted ordered subtree of $\widetilde{\mathscr{T}}^{n}$ spanned by the root and $\{V_{n,1},\dots,V_{n,k}\}$. In this subsection we prove the following proposition on the finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$.

### v2

@@ 4.4.2 Key propositions
By (3.15) in Corollary 3.6,
$\quad\,\mathbb{P}_{n}\Big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\,\Big|\,\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m})\text{ are distinct}\Big)=1-\dfrac{m\sigma_{n}^{2}(j-h)}{q_{n}(j-h)}\big(1+o_{n}(1)\big).$
Substituting into (4.23), we obtain
@@
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
@@
$\quad\cdot\mathbb{P}_{n}\big(\mathfrak{p}(u^{j-h}_{1}),\ldots,\mathfrak{p}(u^{j-h}_{m+1})\text{ are distinct}\big)\cdot\mathbb{P}_{n}\big(\#\big\{\text{Anc}(u^{j-h-1}_{j},h_{r}):1\leq i\leq m\big\}=m\big).$
@@
This proves the lemma. ∎
Proof of Proposition 4.14 (2).
@@
Fix vertices $v^{j}_{l},\,v^{j^{\prime}}_{l^{\prime}}$ such that $h_{r}<j\leq j^{\prime}\leq h_{r-1}$, $1\leq l\leq q_{n}(j)-m$, and $1\leq l^{\prime}\leq q_{n}(j^{\prime})-m$. It suffices to show that, uniformly in $1\leq k^{\prime}\leq j^{\prime}-h_{r}$,
${\mathbb{P}}_{n}\big(\tau(v^{j}_{l})=\infty~\big|~\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big)$ (4.24)
$=$ $(1+o_{n}(1))\prod_{h=0}^{j-(j^{\prime}-k^{\prime})-2}\Bigg[1-\frac{(m+1)\sigma^{2}}{q_{n}(j-h)}(1+o_{n}(1))\Bigg]\cdot\prod_{h={j-(j^{\prime}-k^{\prime})}}^{j-{h_{r}-1}}\Bigg[1-\frac{m\sigma^{2}}{q_{n}(j-h)}(1+o_{n}(1))\Bigg].$
@@
$q_{n}(h):={\mathbb{P}}_{n}\big(\tau(v^{j}_{l})>h+1~\big|~\tau(v^{j}_{l})>h,\,\tau(v^{j^{\prime}}_{l^{\prime}})=k^{\prime},\,v^{j^{\prime}}_{l^{\prime}}\notin\text{Desc}(v^{j}_{l}),\,\widetilde{B}_{r}\big).$
Under this conditioning, $j^{\prime}-k^{\prime}$ represents the height at which $v^{j^{\prime}}_{l^{\prime}}$ merges into $U_{i,r}$. Taking $h_{r-1}$ as the base height, we obtain by exchangeability that:
@@
    forms an (m,jh,m+2)(m,j-h,m+2)-uniform selection pair.
@@
    forms an (m,jh,m+1)(m,j-h,m+1)-uniform selection pair.
@@
    Uniformly over $0\leq h\leq j-h_{r}-1$ with $j-h>j^{\prime}-k^{\prime}+1$,
@@
    Uniformly over $0\leq h\leq j-h_{r}-1$ with $j-h\leq j^{\prime}-k^{\prime}$,
@@
    $q_{n}(h)$ $=\mathbb{P}_{n}\Big(\mathfrak{p}(u_{2}^{j-h}),\ldots,\mathfrak{p}(u_{m+2}^{j-h})\text{ are distinct}~\Big|~\mathfrak{p}(u_{2}^{j-h}),\ldots,\mathfrak{p}(u_{m+1}^{j-h})\text{ are distinct},\,\mathfrak{p}(u_{1}^{j-h})=\mathfrak{p}(u_{2}^{j-h})\Big)$
@@
This completes the proof. ∎
@@
In the previous section we established the tightness of $\widetilde{\mathscr{T}}^{n}$. Here we first prove the finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$ (Section 5.1); combined with tightness, this yields convergence of the contour functions. We then show in Section 5.2 that the discrepancy between the height and contour functions vanishes in the limit, so both processes converge to the same limiting process, which completes the proof of Theorem 1.1.
@@ 5.1 Finite-dimensional convergence of $\widetilde{\mathscr{T}}^{n}$
[…]